7. [-14 Points] DETAILS LARLINALG8 7.1.019. Find the characteristic equation and the eigenvalues and corresponding eigenvectoplease solve both
Answer #1

제 Let A : 1 -) 끼 characteristic eq eqa i 승 | 4 - | = 0. 1 - 기 늘 0. - 2. 세 دیاکے 20, + 2 U =0 (1)(1-A) + = - (1-1)(1+1)+ - [-츠 For engen male solve, 2. 0 - ㅗ 9 =4 SO (1,4> - (2) Let X, be esgeuvector of A Correspondeus to da 2. then 0 (44) = AVI- + 050 :11 - Let 지 Xz be = eugenvector of a comesponding to 1=1 2 then (4-1/2) x2 = 0. 23 늘 JO -3 -3 X2 co rya 2 +H₂O x=-X₂ put X